OpenStudy (anonymous):

Show how the double angle formula for cosine is derived from the sum formula?

OpenStudy (anonymous):

well:\[\cos(2x)=\cos(x+x)\]

OpenStudy (anonymous):

and then the sum formula is used on x+x

OpenStudy (anonymous):

actually it should be:\[\cos^2x-\sin^2x\]

OpenStudy (anonymous):

\[\cos(x+x)=\cos(x)*\cos(x)-\sin(x)*\sin(x)\]

OpenStudy (anonymous):

Ok